An object’s position is described by the equation x=At-Bt^4+C . What are the dimensions of (a)A (b)B (c)C (d) Find the derivative of position dx/dt (e) What are the dimensions of the time derivative?


1
Expert's answer
2021-02-18T18:53:50-0500

a)


"[A]=LT^{-1}"

b)


"[B]=LT^{-4}"

c)


"[C]=L"

d)


"\frac{dx}{dt}=A-4Bt^3"

e)


"\left[\frac{dx}{dt}\right]=LT^{-1}"